OpenStudy (anonymous):

can someone help me with operations of radicals? I will fan and/or medal. depends on how many reply

OpenStudy (anonymous):

Simplify.7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54 11 square root of 6 end root minus 7 square root of 12 11 square root of 3 end root minus 7 square root of 6 negative 3 square root of 9 4 square root of 9

OpenStudy (camerondoherty):

Is this ur Equation?\[7\sqrt{3}- 4\sqrt{6}+ \sqrt{40}-\sqrt{54}\]

OpenStudy (mokeira):

your question is not clear

OpenStudy (anonymous):

yes tht is the question @camerondoherty

OpenStudy (dan815):

\[7\sqrt{3}- 4\sqrt{6}+ \sqrt{48}-\sqrt{54}\]

OpenStudy (camerondoherty):

oh... whoopsies c:

OpenStudy (anonymous):

oh yeah 48

OpenStudy (dan815):

ok carry on

OpenStudy (anonymous):

\[7 \sqrt{3}-4 \sqrt{6}+\sqrt{48}-\sqrt{54}=11 \sqrt{3}-7 \sqrt{6} \]b., the second one.

OpenStudy (anonymous):

but could u explain me how to do it? I want to do it alone in the future

OpenStudy (anonymous):

\[7 \sqrt{3}-4 \sqrt{6}+\sqrt{48}-\sqrt{54}=7 \sqrt{3}-4 \sqrt{6}+4 \sqrt{3}-3 \sqrt{6} \]

OpenStudy (anonymous):

Add up the tems with common radical values.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

thanks for ur help